OpenStudy (anonymous):

Please Help with very quick question! find range of f(x)=-x+4 for the domain -3,-2,-1,1

OpenStudy (anonymous):

Answer choices

OpenStudy (danjs):

The domain is all the possible values for the independent variable, X in this case

OpenStudy (danjs):

The range is all the Y values corresponding to the domain.

OpenStudy (danjs):



OpenStudy (danjs):

f(x) = x - 4 Example f(-3) = (-3) - 4 = -7

OpenStudy (danjs):

Range = {-7 , f(-2), f(-1) , f(1) }

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

f(x)=-x+4 for the domain 4,3,2,0 would be f(0)=-(0)+4=4 f(0)=-(2)+4=2 f(0)=-(3)+4=1 f(0)=-(4)+4=0 the answer would be { 4, 2, 1, 0} you need: f(x)=-x+4 for the domain -3,-2,-1,1 find the range, as I did in my example

OpenStudy (anonymous):

so the range is 1.3 @SolomonZelman

OpenStudy (solomonzelman):

you will have for numbers for the range

OpenStudy (solomonzelman):

*4 numbers

OpenStudy (solomonzelman):

f(x)=-x+4 for the domain -3,-2,-1,1 would be f(0)=-(-3)+4=? f(0)=-(-2)+4=? f(0)=-(-1)+4=? f(0)=-( 1 )+4=? the answer would be { __ , __ , __ , __}

OpenStudy (anonymous):

@SolomonZelman what would be in the zeros??

OpenStudy (solomonzelman):

my bad

OpenStudy (solomonzelman):

it should be saying: (x)=-x+4 for the domain -3,-2,-1,1 would be f(-3)=-(-3)+4=? f(-2)=-(-2)+4=? f(-1)=-(-1)+4=? f( 1 )=-( 1 )+4=? the answer would be { __ , __ , __ , __}

OpenStudy (solomonzelman):

I am plugging each number of range, into the funciton, instead of x. The "output" will be the range

OpenStudy (anonymous):

so 2 is the output

OpenStudy (anonymous):

IM so confused! :'(

OpenStudy (solomonzelman):

answer 4 questions: f(-3)=-(-3)+4=? f(-2)=-(-2)+4=? f(-1)=-(-1)+4=? f( 1 )=-( 1 )+4=?

OpenStudy (anonymous):

I got it!!!!!!!!!!!!!!!!!!!!!!! Bless your heart! Thanks! @SolomonZelman

OpenStudy (solomonzelman):

sure, so what is rthe answer you got?

OpenStudy (solomonzelman):

and tnx for the blessing, loL